OpenStudy (abb0t):

How do I find the particular solution to: y''-2y'-3y=3-10sin(t) I already know the complimentary solution, y(t) = c1e^-t+c2e^3t I just need to know how to solve the particular using 3-10sin(t). Any help would be great assistance.

OpenStudy (anonymous):

Hmmm, it's been a while since I've done these, but my guess is that the particular solution might be of the form \[ A\cos(t)+B\sin(t)+C \]

OpenStudy (abb0t):

Oh yeah, I forgot ot mention I got that, but I am having toruble after i find the derivatives. Because then i have cos and sin. how do I do it so that I get systems. like A+B = # and so forth..

OpenStudy (anonymous):

Well, find y' and y''.

OpenStudy (abb0t):

y = At + Bsin(t) + Dcos(t) y' = A + Bcos(t) - Dsin(t) y' = -Bsin(t) - Dcos(t) -Bsin(t)-Dcos(t)-2(A+Bcos(t)-Dsin(t)) - 3(At+Bsin(t)+Dcos(t)) = 3 - 10sin(t) is that correct so far?

OpenStudy (anonymous):

Yeah, sorry was distracted. I think you should try solving for A, B, and C.

OpenStudy (anonymous):

somehow C became D though

OpenStudy (abb0t):

Yeah, I used D instead of C. That's where i'm stuck on, how do I solve for A,B, and "C".

OpenStudy (anonymous):

First thing you want to do is just simplify the thing.

OpenStudy (anonymous):

You want it in the form of \[ (A + C + blah)\sin(t)+(blah_2)\cos(t)+blah_3 \] or something like that.

OpenStudy (abb0t):

Oh!!! Ok. I see it now! Thanks for the help!! I was trying to separate it in some form but I didn't know how. Thanks!!